# Evan Munro    Xu Kuang    Stefan Wager

### v3

@@ B.11 Proof of Theorem 13
For convenience, we use the following short-hand for the policy value and the asymptotic budget imbalance:
$V(\nu)=\mathbb{E}\left[\nu(X_{i})\tau_{\text{CADE}}^{*}(X_{i})\right],\ \ \ \ G(\nu)=\mathbb{E}\left[\left(\nu(X_{i})-\pi(X_{i})\right)\tau_{\text{CADE}}^{*,z}(X_{i})\right].$ (66)
Then, following the setup of Luenberger (1969, Chapter 8.6), we define primal and dual objectives over $\theta,\,\lambda\in\mathbb{R}_{+}^{J}$ as
$\omega(\theta)=\inf_{\nu}\left\{-V(\nu):\left\lvert G_{j}(\nu)\right\rvert\leq\theta_{j},\,j=1,\,\ldots,\,J\right\},\ \ \ \ \phi(\lambda)=\inf_{\nu}\left\{-V(\nu)+\lambda\cdot\left\lvert G(\nu)\right\rvert\right\},$ (67)
@@
$\omega(\theta)=\sup\left\{\phi(\lambda):\lambda\geq\theta\right\}$ (68)
for all $\theta$ with $\min\left\{\theta_{j}\right\}>0$. The dual function $\phi$ is clearly continuous at 0, and so we also see that for any sequence $\theta$ converging to 0,
$\lim_{\theta\rightarrow 0,\,\theta>0}\omega(\theta)=\sup\left\{\phi(\lambda):\lambda\geq 0\right\}.$ (69)
At the end of this proof, we will further show that $\omega(\theta)$ is continuous at 0, thus implying that in fact
$\omega(0)=\sup\left\{\phi(\lambda):\lambda\geq 0\right\},$ (70)
where $\omega(0)$ is the negative of the optimal value in our original optimization problem.
Now, given strong duality as in (70), we can express the policy value in our original problem as
$\begin{split}V(\nu^{*})&=\inf_{\lambda\in\mathbb{R}^{J}_{+}}\left\{\sup_{\nu}\left\{V(\nu)-\lambda\cdot\left\lvert G(\nu)\right\rvert\right\}\right\}.\end{split}$ (71)
Writing $\nu^{*}_{\lambda}$ for the solution of the inner optimization problem we see immediately from its first-order conditions that, for all $x$, one of the following 3 conditions must hold:
$\tau^{*}_{\text{CADE}}(x)-\sum_{j=1}^{J}\lambda_{j}\operatorname{sign}\left(G_{j}(\nu_{\lambda}^{*})\right)\tau^{*,z}_{CADE,j}(x)>0\ \text{ and}\ \nu_{\lambda}^{*}(x)=1,$
$\tau^{*}_{\text{CADE}}(x)-\sum_{j=1}^{J}\lambda_{j}\operatorname{sign}\left(G_{j}(\nu_{\lambda}^{*})\right)\tau^{*,z}_{CADE,j}(x)<0\ \text{ and}\ \nu_{\lambda}^{*}(x)=0,$
$\tau^{*}_{\text{CADE}}(x)-\sum_{j=1}^{J}\lambda_{j}\operatorname{sign}\left(G_{j}(\nu_{\lambda}^{*})\right)\tau^{*,z}_{CADE,j}(x)=0.$
Thus, equivalently, using notation from (37) we must have
@@
Plugging back into (71), we get
@@
Write $c^{*},\,a^{*}$ for the solution of the above optimization problem. We have
$V(\nu^{*})=V(\nu_{c^{*},a^{*}}(x))-\left\lvert c^{*}\right\rvert\cdot\left\lvert G(\nu_{c^{*},a^{*}}(x))\right\rvert\leq V(\nu_{c^{*},a^{*}}(x)),$
and so $V(\nu^{*})=V(\nu_{c^{*},a^{*}}(x))$ by optimality of the policy $\nu^{*}$.
We now return to verifying continuity of $\omega(\theta)$ at 0. Let
=$\mathcal{B}=\left\{b:\mathbb{E}\left[\left(\nu(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]=b\text{ for some }0\leq\nu(\cdot)\leq 1\right\}.$
[…]
@@
For simplicity, [^7] If the set doesn’t span $\mathbb{R}^{J}$ we can simply ignore the direction without any variation at the cost of extra notation. assume that $\mathcal{B}$ spans $\mathbb{R}^{J}$; thus, by (72) and convexity of $\mathcal{B}$, we know that there is a $\varepsilon>0$ such that $b\in\mathcal{B}$ for all $\left\lVert b\right\rVert_{2}\leq\varepsilon$.
Now, let $\omega(\theta)$ denote the primal objective (67) for some $\left\lVert\theta\right\rVert_{2}\leq\varepsilon$, and let $\nu^{*}_{\theta}$ be a maximizer of this objective. By the above argument, we know that there exists another policy $\nu^{\prime}_{\theta}$ with
$\mathbb{E}\left[\left(\nu^{\prime}_{\theta}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]=-\varepsilon\ \frac{\mathbb{E}\left[\left(\nu^{*}_{\theta}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]}{\left\lVert\mathbb{E}\left[\left(\nu^{*}_{\theta}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]\right\rVert_{2}}.$
We can construct a third policy by convex averaging of $\nu^{*}_{\theta}$ and $\nu^{\prime}_{\theta}$:
$\nu^{+}_{\theta}=\frac{\left\lVert\mathbb{E}\left[\left(\nu^{*}_{\theta}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]\right\rVert_{2}\nu^{\prime}_{\theta}+\varepsilon\nu^{*}_{\theta}}{\left\lVert\mathbb{E}\left[\left(\nu^{*}_{\theta}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]\right\rVert_{2}+\varepsilon}.$
Clearly this policy exactly satisfies the budget constraint (i.e., is a feasible solution to the original optimization problem), and by boundedness of outcomes
$V(\nu^{+}_{\theta})=V(\nu^{*}_{\theta})+\mathcal{O}\left(\left\lVert\theta\right\rVert_{2}/\varepsilon\right).$
Now, because $\nu^{+}_{\theta}$ is feasible we have
$\omega(0)\leq-V(\nu^{+}_{\theta})=-V(\nu^{*}_{\theta})+\mathcal{O}\left(\left\lVert\theta\right\rVert_{2}\right)=\omega(\theta)+\mathcal{O}\left(\left\lVert\theta\right\rVert_{2}\right),$
which implies continuity of $\omega$ at 0 because, by construction, $\omega$ is decreasing in $\theta$.

### v4

@@ B.11 Proof of Theorem 13
and the constraint functions $G_{j}(\nu)=\mathbb{E}\left[\left(\nu(X_{i})-\pi(X_{i})\right)\tau_{\text{CADE},j}^{*,z}(X_{i})\right]$ can be re-expressed as
@@
Now, let
=$\mathcal{B}=\left\{b:G(\nu)=b\text{ for some }0\leq\nu(\cdot)\leq 1\right\},$ (75)
and let $\operatorname{span}\left(\mathcal{B}\right)\subseteq\mathbb{R}^{J}$ be its linear span. Then, Theorem 3.1 of Dantzig & Wald (1951) tells us that, as long as 00 is an interior point of $\mathcal{B}$ with respect to $\operatorname{span}\left(\mathcal{B}\right)$, then the supremum of (73) subject to (74) is achieved, and any $S_{\nu}$ satisfying the constraints (74) is value-maximizing if and only there exists $c\in\mathbb{R}^{J}$ such that, almost surely
$\begin{split}&\tau_{\text{CADE}}^{*}(X_{i})\geq\sum_{j=1}^{J}c_{j}\tau_{\text{CADE},j}^{*,z}(X_{i})\text{ whenever }A\in S_{\nu},\\ &\tau_{\text{CADE}}^{*}(X_{i})\leq\sum_{j=1}^{J}c_{j}\tau_{\text{CADE},j}^{*,z}(X_{i})\text{ whenever }A\not\in S_{\nu}.\end{split}$ (76)
Now, since $U_{i}$ is drawn uniformly over $[0,1]$, a policy $\nu$ induces a set $S_{\nu}$ satisfying (76) if and only if, for each $x\in\mathcal{X}$ except on a set of measure zero:
  1. 1.
    $\tau_{\text{CADE}}^{*}(x)>c^{\top}\tau_{\text{CADE}}^{*,z}(x)$ and $\nu(x)=1$,
  2. 2.
    $\tau_{\text{CADE}}^{*}(x)<c^{\top}\tau_{\text{CADE}}^{*,z}(x)$ and $\nu(x)=0$,
  3. 3.
    $\tau_{\text{CADE}}^{*}(x)=c^{\top}\tau_{\text{CADE}}^{*,z}(x)$ and $\nu(x)\in[0,1]$.
Thus, any policy is optimal if and only if it satisfies the constraints (74) and is (almost surely) of the above form.
It remains to determine behavior of $\nu$ in the region where $\tau_{\text{CADE}}^{*}(x)=c^{\top}\tau_{\text{CADE}}^{*,z}(x)$. Given $c$, for any policy of the above form, satisfying the constraints (74) entails
$\begin{split}&c^{\top}\mathbb{E}\left[\left(\nu(X_{i})-\pi(X_{i})\right)\tau_{\text{CADE}}^{*,z}(X_{i})1\left(\left\{\tau_{\text{CADE}}^{*}(x)=c^{\top}\tau_{\text{CADE}}^{*,z}(x)\right\}\right)\right]\\ &\quad\quad+c^{\top}\mathbb{E}\left[\left(1-\pi(X_{i})\right)\tau_{\text{CADE}}^{*,z}(X_{i})1\left(\left\{\tau_{\text{CADE}}^{*}(x)>c^{\top}\tau_{\text{CADE}}^{*,z}(x)\right\}\right)\right]\\ &\quad\quad\quad\quad+c^{\top}\mathbb{E}\left[\left(0-\pi(X_{i})\right)\tau_{\text{CADE}}^{*,z}(X_{i})1\left(\left\{\tau_{\text{CADE}}^{*}(x)<c^{\top}\tau_{\text{CADE}}^{*,z}(x)\right\}\right)\right]=0.\end{split}$ (77)
Now, in the 3rd region we’re considering $\tau_{\text{CADE}}^{*}(x)=c^{\top}\tau_{\text{CADE}}^{*,z}(x)$, and so any policy satisfying these constraints must in fact have the same value. Thus, if there exists a constant $b$ such that a policy of the above form with $\nu(x)=b$ for all $x$ $\tau_{\text{CADE}}^{*}(x)=c^{\top}\tau_{\text{CADE}}^{*,z}(x)$ satisfies the constraints, then this policy is in fact optimal.
It now remains to check the interior-point condition. Recall that, in our Design 1, we have “overlap” in the sense that $\eta\leq\pi(x)\leq 1-\eta$ for all $x\in\mathcal{X}$ for some $\eta>0$. The main step is to show that, under our conditions,
@@
To verify this, pick any $b\in\mathcal{B}$. There must be a $\nu(\cdot)$ such that $\mathbb{E}\left[\left(\nu(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]=b$. Given such a $\nu$, let $\tilde{\nu}(\cdot)=\pi(\cdot)-\eta(\nu(\cdot)-\pi(\cdot))$. By overlap, $0\leq\tilde{\nu}(\cdot)\leq 1$, i.e., $\tilde{\nu}(\cdot)$ is a valid policy. Furthermore, clearly, $\mathbb{E}\left[\left(\tilde{\nu}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]=-\eta b$, i.e., $-\eta b\in\mathcal{B}$. Finally, because $\mathcal{B}$ is also convex, (78) immediately implies the interior-point condition: if we pick a maximal set of linearly independent vectors in $\mathcal{B}$, then their mirror images (up to a factor $\eta$) are also in $\mathcal{B}$, and the convex hull of both the original vectors and their mirror images contains an open set around 0 and is in turn contained in $\mathcal{B}$.
B.12 Consistency of Estimated Targeting Rule
When the optimal rule is deterministic and unique, it has the form $\nu^{*}(X_{i})=\mathbbm{1}(\tau^{*}_{\text{CADE}}(X_{i})\geq c^{*}\tau^{*,z}_{\text{CADE}}(X_{i}))$. By Theorem 13, we can write the optimal rule as the unique solution to the following $J$-dimensional moment condition, where $\tau^{*}$ represents the vector of functions that concatenates $\tau^{*}_{\text{CADE}}(X_{i})$ and $\tau^{*,z}_{\text{CADE}}(X_{i})$ and $\tau$ is some bounded function with the same domain and codomain.
$G(c^{*};\tau^{*})=0,\mbox{ where }G(c;\tau)=\mathbb{E}\left[\Big{(}\mathbbm{1}(\tau(X_{i})\geq c^{\top}\tau^{z}(X_{i}))-\pi(X_{i})\Big{)}\tau^{z}(X_{i})\right]$
$c^{*}$ is finite, so we can specify a ball in $\mathbb{R}^{J}$ such that the distance from $c^{*}$ to the surface of this ball is at least $M$. Specify such a ball as $\mathcal{B}_{M}$. We can estimate the optimal rule by solving the following $J$-dimensional score condition. The conditional average direct effects are nuisance functions, since they need to be estimated, where $\hat{\tau}$ is some estimate of the population conditional average treatment effects on outcomes and net demand:
$\begin{split}&\hat{c}\in\{c\in\mathbb{R}^{J}:\hat{G}_{n}(c;\hat{\tau})=o_{p}(1)\},\\ &\hat{G}_{n}(c;\hat{\tau})=\frac{1}{n}\sum\limits_{i=1}^{n}(\mathbbm{1}(\hat{\tau}(X_{i})\geq c^{\top}\hat{\tau}^{z}(X_{i})-\pi(X_{i}))\hat{\tau}^{z}(X_{i}).\end{split}$
The proof of Proposition 18 shows that the optimal-equilibrium neutral rule is the unique (and well-separated) solution to the $J$ population constraints defined by $G(\cdot)$. Then, the consistency of the estimated rule follows from uniform convergence of the empirical constraint functions to the population constraint functions.
Proposition 18.
For every $v\in S^{J-1}$ and $b\in\mathbb{R}^{+}$, where $S^{J-1}$ is the unit sphere in $\mathbb{R}^{J}$, assume that the random variable $\tau^{*}_{\text{CADE}}(X_{i})-b\cdot v^{\top}\tau^{z,*}_{\text{CADE}}(X_{i})$ has uniformly bounded density, and for every $v\in S^{J-1}$ that $v^{\top}\tau^{z,*}_{\text{CADE}}(X_{i})$ has uniformly bounded density. In addition, assume that the covariance matrix of $\tau^{*,z}_{\text{CADE}}(X_{i})$ is positive definite, and that $\hat{G}_{n}(\hat{c};\hat{\tau})=o_{p}(1)$. Finally, assume that $\hat{\tau}:\mathcal{X}\to\mathbb{R}$ and $\hat{\tau}^{z}:\mathcal{X}\to\mathbb{R}$ are functions that are uniformly bounded and are consistent estimators of the population conditional average treatment effects in the following sense:
$\begin{split}&\mathbb{E}_{T}[(\hat{\tau}(X_{i})-\tau^{*}_{\text{CADE}}(X_{i}))^{2}]=o_{p}(1),\\ &\mathbb{E}_{T}[||\hat{\tau}^{z}_{j}(X_{i})-\tau^{*,z}_{j,\text{CADE}}(X_{i}))||_{2}^{2}]=o_{p}(1),\end{split}$
where $\mathbb{E}_{T}[\cdot]$ is the expectation over a random sample of test data, conditional on the training data used for estimation.
Then, the estimated targeting rule is consistent for the optimal equilibrium-stable targeting rule in the population,
$\hat{c}=c^{*}+o_{p}(1).$
First, we use a change of variables to make the parameter space compact. For every $c\in\mathbb{R}^{J}$, we can write $c=f(a)\cdot v$, where $v=c/||c||_{2}\in S^{J-1}$ and $a=f^{-1}(||c||_{2})\in[0,1)$ for some continuous and strictly monotonic function $f:[0,1]\to[0,+\infty]$ with $f(0)=0$ and $f(1)=\infty$. For this proof, it will be useful to characterize properties of the population constraint function over the extended (and compact) space of $a\in[0,1]$, even though the optimal rule, and all of the estimated rules are found within $a\in[0,1)$. For a more compact notation define $\theta$ as the vector combining $v$ and $a$, where $\theta\in\Theta$, and $\Theta=S^{J-1}\times[0,1]$. Let
$\nu(X_{i};\theta,\tau)=\begin{cases}\mathbbm{1}(\tau(X_{i})>f(a)v^{\top}\tau^{z}(X_{i})),&0\leq a<1,\\ \mathbbm{1}(0>v^{\top}\tau^{z}(X_{i})),&a=1.\end{cases}$
Now we can define the population and empirical constraint at a given treatment rule and for a given set of conditional mean functions:
$G(\theta;\tau)=\mathbb{E}_{T}\left[\Big{(}\nu(X_{i};\theta,\tau)-\pi(X_{i})\Big{)}\tau^{z}(X_{i})\right],\qquad G_{n}(\theta^{*};\tau)=\frac{1}{n}\sum\limits_{i=1}^{n}\Big{(}\nu(X_{i};\theta,\tau)-\pi(X_{i})\Big{)}\tau^{z}(X_{i}).$
Any estimated treatment rule, characterized by $\hat{\theta}$, approximately satisfies the empirical constraint with conditional average treatment effects estimated by data-splitting, so that $G_{n}(\hat{\theta};\hat{\tau})=o_{p}(1)$.
Under our assumptions, Theorem 13 indicates that there is a unique and deterministic rule in the population characterized by $G(\theta^{*};\tau^{*})=0$, where $\tau^{*}$ collects $\tau^{*}_{\text{CADE}}(X_{i})$ and $\tau^{*,z}_{\text{CADE}}(X_{i})$ and $\theta^{*}$ collects $a^{*}$ and $v^{*}$. It is deterministic because the set $\{x\in\mathcal{X}:\tau^{*}_{\text{CADE}}(x)=f(a^{*})\cdot v^{*\top}\tau^{z,*}_{\text{CADE}}(x)\}$ has measure 0. By Theorem 13, any optimal equilibrum-neutral rule that is deterministic has the structure $\mathbbm{1}(\tau^{*}_{\text{CADE}}(x)>f(a^{*})v^{*,\top}\tau^{z,*}_{\text{CADE}}(x))$. Since the covariance matrix of $\tau^{*,z}_{\text{CADE}}(X_{i})$ is positive definite, there can be only one such rule that meets the population constraints (this rules out label swapping between two goods in the market, for example).
Next, we want to show that each element of $G(\theta;\tau^{*})$ is a continuous function in $\theta$ for all $\theta\in\Theta$. Choose some $a_{0},v_{0}$ and arbitrary $j\in\{1,\ldots,J\}$. For some sequence $\theta_{n}$, where $\theta_{n}\to\theta_{0}=(a_{0},v_{0})$, by the boundedness of $\tau^{z}(X_{i})$, there is some finite $M$ such that
$\begin{split}|G_{j}(\theta_{n};\tau^{*})-G_{j}(\theta_{0};\tau^{*})|&\leq P(\nu(X_{i};\theta_{n},\tau^{*})\neq\nu(X_{i};\theta_{0},\tau^{*}))M\\ &=O(||\theta_{n}-\theta_{0}||_{2})\\ &=o(1)\end{split}$
The first step is by Lemma 19. This proves that when $a=1$, then $G(\theta;\tau^{*})$ is continuous in $\theta$. Since $G(\theta;\tau)$ is a continuous function in $\theta$, $\hat{\theta}$ is also in a compact space, and $\theta^{*}$ is the unique solution to $G(\theta^{*};\tau^{*})=0$, then we know that if $\epsilon>0$, then $\sup\limits_{\theta:||\theta-\theta^{*}||_{2}>\epsilon}||G(\theta)||_{2}>0$. The final step is to show uniform consistency:
$\sup\limits_{\theta\in\Theta}|\hat{G}_{n}(\theta;\hat{\tau})-G(\theta;\tau^{*})|=o_{p}(1).$
It will be useful to make the following decomposition:
$|\hat{G}_{n}(\theta;\hat{\tau})-g(\theta;\tau)|\leq|G_{n}(\theta;\hat{\tau})-G(\theta;\hat{\tau})|+|G(\theta;\hat{\tau})-G(\theta;\tau^{*})|.$ (79)
For the first term of (79), we write the empirical average using data-splitting, where $I_{k}$ are the indexes of data in split $k$, and $\hat{\tau}^{k(i)}(X_{i})$ are conditional average treatment effects estimated on data that is not in $I_{k}$. We drop the CADE subscripts here to keep the notation more manageable. Then, by treating $\hat{\tau}^{k}$ as fixed within each split, we use tail bounds for empirical processes indexed by VC-classes. The details are as follows:
$\begin{split}&\sup_{\theta\in\Theta}|G_{n}(\theta;\hat{\tau})-G(\theta;\hat{\tau})|\\ &=\sup\limits_{\theta\in\Theta}\left|\frac{1}{n}\sum\limits_{i=1}^{n}(\nu(X_{i};\theta,\hat{\tau})-\pi(X_{i}))\hat{\tau}^{z}(X_{i})-\mathbb{E}_{T}[(\nu(X_{i};\theta,\hat{\tau})-\pi(X_{i}))\hat{\tau}^{z}(X_{i})]\right|\\ &=\sup\limits_{\theta\in\Theta}\left|\sum\limits_{k=1}^{K}n_{k}/n\frac{1}{n_{k}}\sum\limits_{i\in I_{k}}(\nu(X_{i};\theta,\hat{\tau}^{k(i)})-\pi(X_{i}))\hat{\tau}^{k(i),z}(X_{i})-\mathbb{E}_{T}[(\nu(X_{i};\theta,\hat{\tau}^{k(i)})-\pi(X_{i}))\hat{\tau}^{k(i),z}(X_{i})]\right|\\ &\leq\sum\limits_{i=1}^{K}\frac{n_{k}}{n}\sup\limits_{\theta\in\Theta}|A_{nk}(\theta)|\\ &\stackrel{{\scriptstyle(1)}}{{=}}o_{p}(1)\end{split}$
Within each split, we can treat $\hat{\tau}^{k}$ as fixed. =$\mathcal{F}=\{\mathbbm{1}(\tau(X_{i})\geq f(a)v^{\top}\tau^{z}(X_{i}))\tau^{z}(X_{i}):a\in[0,1],v\in S^{J-1}\}$ for a fixed and uniformly bounded $\tau(X_{i})$ is a VC-class multiplied by a bounded random variable. Then, the tail bound for empirical processes indexed by VC-classes (e.g. Theorem 2.6.7 of van der Vaart & Wellner (1996)) gives a tail bound for $\sup_{\theta\in\Theta}|A_{nk}(\theta)|$ that does not depend on a given realization of $\hat{\tau}$. This implies $\lim\limits_{n\to\infty}Pr(\sup\limits_{\theta\in\Theta}|A_{nk}|>t)=0$ for any $t>0$.
To handle the second term in (79),
$\begin{split}\sup_{\theta\in\Theta}|G(\theta;\hat{\tau})-G(\theta;\tau^{*})|&=\sup_{\theta\in\Theta}|\mathbb{E}_{T}[(\nu(X_{i};\theta,\hat{\tau})-\pi(X_{i}))\hat{\tau}^{z}(X_{i})]-\mathbb{E}_{T}[(\nu(X_{i};\theta,\tau^{*})-\pi(X_{i}))\tau^{*,z}(X_{i})]|\\ &\leq\sqrt{\mathbb{E}_{T}[(\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i}))^{2}]}+\sup_{\theta\in\Theta}M\sqrt{\mathbb{E}_{T}[(\nu(X_{i};\theta,\hat{\tau})-\nu(X_{i};\theta,\tau^{*}))^{2}]}\\ &\leq o_{p}(1)+\sup_{\theta\in\Theta}M\sqrt{P(\nu(X_{i};\theta,\hat{\tau})\neq\nu(X_{i};\theta,\tau^{*}))}\\ &\stackrel{{\scriptstyle(1)}}{{=}}o_{p}(1)+O(||\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i})||_{2}+|\tau^{*}(X_{i})-\hat{\tau}(X_{i})|)\\ &=o_{p}(1)\end{split}$
where $M$ is finite since net demand is bounded, (1) is from Lemma 19 and the final step is from the assumption on the mean-square convergence of the nuisance functions. We have now shown that $\sup\limits_{\theta\in\Theta}|\hat{G}_{n}(\theta;\hat{\tau})-G(\theta;\tau^{*})|=o_{p}(1).$
Now, that we have proved that the unique population treatment rule is well-separated and that the sample constraints converge uniformly to the population constraints, following Theorem 5.9 of van der Vaart (1998), we have consistency of $\hat{\theta}$ to $\theta^{*}$. The details follow. Since $\hat{\theta}$ satisfies the empirical constraints, then $||G_{n}(\hat{\theta};\hat{\tau})||_{2}\leq||G_{n}(\theta^{*};\tau^{*})||_{2}+o_{p}(1)$. By the convergence of the empirical constraints in the RHS, $||G_{n}(\hat{\theta};\hat{\tau})||_{2}\leq||G(\theta^{*};\tau^{*})||_{2}+o_{p}(1)$. Adding and subtracting $||G(\hat{\theta};\tau^{*})||_{2}$, we have:
$\begin{split}||G(\hat{\theta};\tau^{*})||_{2}-||G_{n}(\hat{\theta};\hat{\tau})||_{2}&\geq||G(\hat{\theta};\tau^{*})||_{2}-||G(\theta^{*};\tau^{*})||_{2}-o_{p}(1)\\ ||G(\hat{\theta};\tau^{*})||_{2}-||G(\theta^{*};\tau^{*})||_{2}&\leq\sup_{\theta\in\Theta}||G(\theta;\tau^{*})||_{2}-||G_{n}(\theta;\hat{\tau})||_{2}\\ ||G(\hat{\theta};\tau^{*})||_{2}&\leq\sup_{\theta\in\Theta}||G(\theta;\tau^{*})-G_{n}(\theta;\hat{\tau})||_{2}\\ ||G(\hat{\theta};\tau^{*})||_{2}&\leq o_{p}(1)\end{split}$
Since $\theta^{*}$ is well-separated, then for any $\epsilon>0$, for any $\theta$ such that $||\theta-\theta^{*}||>\epsilon$, then $||G(\theta)||_{2}>\eta$. So, for any $\epsilon>0$, $P(||\hat{\theta}_{n}-\theta^{*}||>\epsilon)\leq P(||G(\hat{\theta};\tau^{*})||_{2}>\eta)\rightarrow_{p}0$. Since $\hat{\theta}$ is just a re-parameterization of $\hat{c}$, then we have shown that $\hat{c}=c^{*}+o_{p}(1)$.
Lemma 19.
Under the Assumptions (and notation) of Proposition 18,
$\begin{split}&P(\nu(X_{i};\theta_{n},\tau^{*})\neq\nu(X_{i};\theta_{0},\tau^{*}))=O(||\theta_{n}-\theta_{0}||_{2}),\\ &\sup\limits_{\theta\in\Theta}P(\nu(X_{i};\theta,\hat{\tau})\neq\nu(X_{i};\theta,\tau^{*}))=O(||\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i})||_{2}+|\tau^{*}(X_{i})-\hat{\tau}(X_{i})|).\end{split}$
Proof.
For $P(\nu(X_{i};\theta_{n},\tau^{*})\neq\nu(X_{i};\theta_{0},\tau^{*}))$ we divide into two cases. In the first case, $a_{0}<1$. We can choose large enough $n$ (to ensure that $a_{n}\neq 1$) so that
$\begin{split}&P(\nu(X_{i};\theta_{n},\tau^{*})\neq\nu(X_{i};\theta_{0};\tau^{*}))\\ &=P(\min\{f(a_{0})v_{0}^{\top}\tau^{*,z}(X_{i}),f(a_{n})v_{n}^{\top}\tau^{*,z}(X_{i})\}<\tau^{*}(X_{i})<\max\{f(a_{0})v_{0}^{\top}\tau^{*,z}(X_{i}),f(a_{n})v_{n}^{\top}\tau^{*,z}(X_{i})\})\\ &\leq P\Big{(}|\tau^{*}(X_{i})-f(a_{0})v_{0}^{\top}\tau^{*,z}(X_{i})|<|(f(a_{n})v_{n}-f(a_{0})v_{0})\tau^{*,z}(X_{i})|\Big{)}\\ &\leq P\Big{(}|\tau^{*}(X_{i})-f(a_{0})v_{0}^{\top}\tau^{*,z}(X_{i})|<|f(a_{n})v_{n}-f(a_{0})v_{0}|M\Big{)}\\ &\leq C||\theta_{n}-\theta_{0}||_{2},\end{split}$
for finite $C$, where $M$ is finite since net demand is bounded. The last step is because $\tau^{*}(X_{i})-av^{\top}\tau^{*,z}(X_{i})$ has bounded density, so its distribution function is Lipschitz, and $f(\cdot)$ is continuous. Next, for $a_{0}=1$, we can prove something similar. We can choose large enough (to ensure that $a_{n}\neq 0$) so that
$\begin{split}&P(\nu(X_{i};\theta_{n},\tau^{*})\neq\nu(X_{i};\theta_{0};\tau^{*}))\\ &=P(\min\{0,(v_{0}^{\top}-v_{n}^{\top})\tau^{*,z}(X_{i})+\tau^{*}(X_{i})/a_{n}\}<v_{0}^{\top}\tau^{*,z}(X_{i})<\max\{0,(v_{0}^{\top}-v_{n}^{\top})\tau^{*,z}(X_{i})+\tau^{*}(X_{i})/a_{n}\})\\ &\leq P\Big{(}|v_{0}^{\top}\tau^{*,z}(X_{i})|<|(v_{0}^{\top}-v_{n}^{\top})\tau^{*,z}(X_{i})+\tau^{*}(X_{i})/f(a_{n})|)\\ &\leq P\Big{(}|v_{0}^{\top}\tau^{*,z}(X_{i})|<|(v_{0}^{\top}-v_{n}^{\top})M+M/f(a_{n})|)\\ &\leq C||\theta_{n}-\theta_{0}||_{2}\end{split}$
for some finite $C$, where $M$ is finite by boundedness of outcomes and net demand. The last step is because $v^{\top}\tau^{*,z}(X_{i})$ has bounded density, so its distribution function is Lipschitz, and $f(\cdot)$ is continuous.
For the second part of the Lemma, also split into two cases. First, when $0\leq f(a)\leq 2$ (denote the product of this restricted space for $a$ and $S^{J-1}$ as $\Theta^{+}$), let $b_{n}=av^{\top}(\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i}))+\tau^{*}(X_{i})-\hat{\tau}(X_{i})$.
$\begin{split}&\sup_{\theta\in\Theta^{+}}P(\nu(X_{i};\theta,\hat{\tau})\neq\nu(X_{i};\theta,\tau^{*}))\\ &=P(\min\{b_{n},0\}<\tau^{*}(X_{i})-f(a)v^{\top}\tau^{*,z}(X_{i})<\max\{b_{n},0\})\\ &\leq\sup_{\theta\in\Theta^{+}}P(|\tau^{*}(X_{i})-f(a)v^{\top}\tau^{*,z}(X_{i})|\leq|f(a)v^{\top}(\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i}))+\tau^{*}(X_{i})-\hat{\tau}(X_{i})|)\\ &\leq\sup_{\theta\in\Theta^{+}}P(|\tau^{*}(X_{i})-f(a)v^{\top}\tau^{*,z}(X_{i})|\leq 2||\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i})||_{2}+|\tau^{*}(X_{i})-\hat{\tau}(X_{i})|)\\ &=O(||\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i})||_{2}+|\tau^{*}(X_{i})-\hat{\tau}(X_{i})|)\\ \end{split}$
(1) is because $\tau^{*}(X_{i})-av^{\top}\tau^{*,z}(X_{i})$ has a density uniformly bounded over $a$ and $v$. This means that the distribution function of $\tau^{*}(X_{i})-av^{\top}\tau^{*,z}(X_{i})$, which we can call $F(t)$ is Lipschitz in $t$, so
$F(2||\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i})||_{2}+|\tau^{*}(X_{i})-\hat{\tau}(X_{i})|)-F(0)\leq 2M||\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i})||_{2}+M|\tau^{*}(X_{i})-\hat{\tau}(X_{i})|,$
where $M$ does not depend on $a$ or $v$. When $a\geq 2$, we can argue similarly, where we use $\Theta^{-}$ to denote the product of this restricted space and $S^{J-1}$. Let $c_{n}=v^{\top}(\tau^{*,z}(X_{i})-\hat{\tau}^{*,z}(X_{i}))+1/f(a)(\hat{\tau}^{*}(X_{i})-\tau^{*}(X_{i}))$.
$\begin{split}&\sup_{\theta\in\Theta^{-}}P(\nu(X_{i};\theta,\hat{\tau})\neq\nu(X_{i};\theta,\tau^{*}))\\ &=P(\min\{c_{n},0\}<av^{\top}\tau^{*,z}(X_{i})-1/a\tau^{*}(X_{i})<\max\{c_{n},0\})\\ &\leq\sup_{\theta\in\Theta^{-}}P(v^{\top}\tau^{*,z}(X_{i})-(1/f(a))\tau^{*}(X_{i})\leq||\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i})||_{2}+|\tau^{*}(X_{i})-\hat{\tau}(X_{i})|)\\ &\leq O(||\hat{\tau}^{z}(X_{i})-\tau^{*,z}(X_{i})||_{2}+|\tau^{*}(X_{i})-\hat{\tau}(X_{i})|)\\ \end{split}$
Where for the last step this is because under our assumptions, for any $f(a)\in[2,\infty]$ the density of $v^{\top}\tau^{*,z}(X_{i})-1/f(a)\tau^{*}(X_{i})$ is uniformly bounded (the bound does not depend on $v$ or $a$). ∎